# Adapters for Altering LLM Vocabularies: What Languages Benefit the Most?

### v1

@@ Abstract
Vocabulary adaptation, which integrates new vocabulary into pre-trained language models (LMs), enables expansion to new languages and mitigates token over-fragmentation. However, existing approaches are limited by their reliance on heuristic or external embeddings. We propose VocADT, a novel method for vocabulary adaptation using adapter modules that are trained to learn the optimal linear combination of existing embeddings while keeping the model’s weights fixed. VocADT offers a flexible and scalable solution without requiring external resources or language constraints. Across 11 languages—with various scripts, resource availability, and fragmentation—we demonstrate that VocADT outperforms the original Mistral model (Jiang et al., 2023) and other baselines across various multilingual tasks. We find that Latin-script languages and highly fragmented languages benefit the most from vocabulary adaptation. We further fine-tune the adapted model on the generative task of machine translation and find that vocabulary adaptation is still beneficial after fine-tuning and that VocADT is the most effective method. [^1] Project page: https://github.com/h-j-han/VocADT. Models at Huggingface Hub
@@ 1 Introduction
Vocabulary adaptation (or transfer)—a process of modifying a pre-trained language model (LM) to use a new vocabulary—offers several key advantages. First, it enables the introduction of new languages into a model, allowing for greater flexibility in handling linguistic diversity and improving downstream performance in target languages (Wang et al., 2020; Gogoulou et al., 2022; Downey et al., 2023). Second, it addresses the problem of over-fragmentation, where tokens are excessively split, potentially slowing down generation and impairing performance in certain languages (Ahia et al., 2023; Petrov et al., 2023; Yamaguchi et al., 2024). These benefits have led to the development of numerous vocabulary adaptation approaches that initialize the new embeddings of new vocabulary with various methods based on heuristics (Mosin et al., 2023; Gee et al., 2022; Downey et al., 2023), external resources (Tran, 2020; Dobler & de Melo, 2023; Liu et al., 2024), or a separate hypernetwork that generates it (Minixhofer et al., 2024), optionally followed by continued training to finalize the adaptation (Minixhofer et al., 2022; Ostendorff & Rehm, 2023; Dobler & de Melo, 2023; Liu et al., 2024).
@@
However, existing vocabulary adaptation approaches face several limitations. Those that rely on heuristics (Gee et al., 2022; Downey et al., 2023), often lack adaptability and require an additional training phase of full-weight updates to fully adapt to the new vocabulary. Those that depend on external embeddings or networks (Tran, 2020; Dobler & de Melo, 2023; Liu et al., 2024), increase complexity and limit scalability. Furthermore, some approaches focus solely on language-specific cases or may have restrictions on the number of languages in the implementation when configuring the vocabulary (Dobler & de Melo, 2023; Minixhofer et al., 2024).
@@
Additionally, we still know little about the impact of vocabulary adaptation across diverse linguistic and task settings. Most prior work investigates a small number of languages, which is insufficient for identifying which languages benefit most from adaptation (Tran, 2020; Ostendorff & Rehm, 2023; Remy et al., 2023; Yamaguchi et al., 2024), while studies that consider a broader range of languages only report averaged results without detailed analysis (Liu et al., 2024; Mundra et al., 2024). Furthermore, the impact of vocabulary adaptation on cross-lingual and generative tasks like machine translation (MT) is understudied, even though they represent crucial application areas for porting models to new languages. Many adaptation methods (Chung et al., 2020; Gee et al., 2022; Liu et al., 2024) have been evaluated instead on non-cross-lingual and discriminative tasks such as commonsense reasoning, natural language inference (NLI), or question answering (QA)—which are typically classification tasks.
@@
We propose VocADT, a novel solution for vocabulary adaptation using adapters, designed to address key challenges in existing approaches. We introduce a vocabulary adapter module between the new vocabulary and the original embeddings of a language model. The module gradually adapts to new vocabularies through training while keeping all weights of the original model fixed, allowing the module to learn the best combination of the original embeddings. This learned adaptation approach makes it more adaptable to new languages and more flexible in the number of languages while removing the necessity of external pre-trained resources.
@@
In addition to our novel method, we empirically address the following key questions to understand the effectiveness and behavior of vocabulary adaptation: (1) Which languages benefit most from vocabulary adaptation? (2) What are the best strategies for creating new vocabularies; is script consistency necessary? (3) How does vocabulary adaptation impact machine translation? We emphasize this task as it is a critical task for multilingual models that involves cross-lingual and generative capabilities, which are often more complex than classification or monolingual tasks.
We demonstrate the effectiveness of our adaptation method on various NLP tasks spanning Natural Language Understanding and MT. Results show that our approach consistently surpasses the original Mistral model in most cases, both after the adaptation phase and following phase of full-weight training. Additionally, our method outperforms or matches other strong vocabulary adaptation baselines. Our findings indicate that Latin-script languages and those with severe fragmentation benefit the most from vocabulary adaptation, while preserving embeddings of overlapped tokens proves useful for non-Latin languages, but not for Latin languages. Finally, while all vocabulary adaptation methods continue to be effective for machine translation after fine-tuning, VocADT shows the best results among them. Our main contributions are summarized as follows:
  • We propose VocADT, a simple and effective solution for vocabulary adaptation using adapters, that addresses key challenges such as reliance on external embedding or language constraints.
  • We design experiments that cover a wide range of languages and scripts, finding that languages with Latin scripts or severe fragmentation benefit the most and that having a consistent grouping of scripts for multilingual vocabulary is helpful.
  • Our approach consistently outperforms the original language model and is more effective than, or at least on par with, strong vocabulary adaptation baselines after the adaptation phase across various tasks and after the following full-weight fine-tuning on MT.
@@
Figure 1: Overview of our vocabulary adaptation with adapter (VocADT) and the initialization of adapter. The vocabulary adapter modules are trained to adapt new vocabulary with existing embeddings while keeping the original model fixed. We initialize entries of the adapter for overlapping tokens and tokens whose partitions are in the original vocabulary. Once trained, the adapters and original embeddings are combined to form the new embeddings.
2 Background
Approaches
Many vocabulary adaptation methods focus on initialization strategies for the new vocabulary embeddings, before continuing training with unlabeled target language text using the original self-supervised pretraining objective. For instance, FOCUS (Dobler & de Melo, 2023) initializes embeddings as a weighted combination of overlapping tokens using external embeddings for non-overlapping ones, while copying the embeddings of overlapping tokens. OFA (Liu et al., 2024) also relies on external word vectors to initialize embeddings for non-shared new tokens, using a weighted average of original tokens based on semantic similarity. This strategy often requires external resources such as auxiliary embeddings (Dobler & de Melo, 2023; Liu et al., 2024; Ostendorff & Rehm, 2023) or bilingual dictionaries (Mundra et al., 2024; Minixhofer et al., 2022).
After initialization, language adaptive pretraining (LAPT; Chau et al., 2020) typically updates all model weights (Tran, 2020; Liu et al., 2024; Dobler & de Melo, 2023; Ostendorff & Rehm, 2023), except Yamaguchi et al. (2024) which use LoRA (Hu et al., 2022). Downey et al. (2023) show that full-weight updates outperform embedding-only training, which is insufficient for multilingual model transfer.
@@ 2 Background
Furthermore, most studies exclusively evaluate models on non-cross-lingual and non-generative tasks, such as binary or multi-class classification, sequence labeling (e.g., part-of-speech tagging), or answer span prediction. Mundra et al. (2024) and Yamaguchi et al. (2024) include monolingual generative tasks like summarization. As a result, the impact of vocabulary adaption on cross-lingual generation tasks such as MT remains understudied, even though this is a crucial application area.  [^2] Mundra et al. (2024) report MT results, but do not release model, code, and number of individual languages to enable comparisons.
To address these gaps, this work introduces a lightweight strategy to adapt LLM to new vocabularies, and experiments designed to measure its impact across diverse linguistic and task settings. We compare empirically against FOCUS and OFA, representing more resource-intensive initialization approaches, and ZeTT, representing a more parsimonious approach that has only been tested on limited tasks so far. Summaries can be found in Appendix Table 9.
@@ 3.1 Vocabulary Adapter Module
We introduce the vocabulary adapter modules to find parameters of new embeddings that can replace the original embedding without changing the non-embedding part of the original model (Figure 1(a)). For simplicity, we refer to both input and output embeddings (or the LM head) collectively as embeddings. Let $V^{o}$ and $V^{n}$ be the $\bm{o}$riginal and $\bm{n}$ew vocabulary, respectively, and let $\mathcal{T}^{x}:w\rightarrow(t_{1},t_{2},\dots,t_{k})$ be a tokenizer associated with a vocabulary $V^{x}$ where $t_{j}\in V^{x},\forall j=1,\dots,k$. We put vocabulary adapter modules ${\bm{A}}\in\mathbb{R}^{|V^{n}|\times|V^{o}|}$ between the new vocabulary $V^{n}$ and the original embedding ${\bm{E}}^{o}\in\mathbb{R}^{|V^{o}|\times h}$ where $h$ is an embedding dimension, in a manner similar to bottleneck adapters (Houlsby et al., 2019). We train the adapters with the standard language modeling loss $\mathcal{L}^{lm}$, where we freeze the original weights and only update the adapters. This may be analogous to finding the new embedding vector for a token with the weighted combination of original embedding vectors (Downey et al., 2023; Dobler & de Melo, 2023; Liu et al., 2024). Unlike similar works, our approach learns the combination with adapter training. After training the adapters, we get new embeddings ${\bm{E}}^{n}\in\mathbb{R}^{|V^{n}|\times h}$ by merging the original embeddings and adapters to ${\bm{E}}^{n}={\bm{A}}{\bm{E}}^{o}$, which result in a language model with the same architecture as the original one but with a different vocabulary size.
@@ 3.4 Further Fine-tuning For Downstream Task
Does the impact of vocabulary adaptation persist after fine-tuning the vocabulary adapted language model for strong performance in a target task? We address this question for the cross-lingual generation task of machine translation, following the training recipe of ALMA (Xu et al., 2024) after the vocabulary adaptation process. ALMA training begins with fine-tuning on monolingual data, followed by further weight optimization on small high-quality parallel data. This two-step fine-tuning process ensures strong translation performance. During the VocADT adaptation phase, we only update the new embeddings. Subsequently, and then update all model parameters in the task-oriented fine-tuning phases.
@@ 4 Which Languages Benefit the Most from Vocabulary Adaptation?
We aim to understand “When and how should we perform vocabulary adaptation?”. More specifically, we seek insight into which languages might benefit the most from vocabulary adaptation in terms of improving overall performance or mitigating over-fragmentation.333 All non-English that we cover tend to be highly fragmented by a vocabulary of popular large language models such as Mistral (Figure 4), resulting in similar effects of mitigating over-fragmentation. Consequently, we emphasize the analysis of the performance aspect. To this end, we design experiments to cover 10 non-English languages along with English, listed in Table 1, with a variety of scripts and language families. These languages are broadly categorized into three groups. The first one is the Latin group of Swahili, Indonesian, Estonian, and Haitian, which are low- to mid-resource languages and all use Latin script. The second is a Mixed group including Korean, Greek, Russian, and Bulgarian, which utilize a mixture of scripts. We also experiment with grouping the Cyrillic script (Cyrillic group) languages.
@@ 5.2 Training
In the adapter training phase, we train only the adapters, while fixing all parameters of the original model. The input and output adapters are separate modules, as preliminary results showed that sharing an adapter for the input and output sides performs worse. We train 3.5B monolingual tokens mixed by 5 languages (English + 4 non-English from each corresponding group) and report test numbers from it. We use “clean” documents from the corpus of MADLAD-400 (Kudugunta et al., 2023). We set the weighing factor of auxiliary loss $\alpha$ with 0.1 for non-Latin groups and 0 for the Latin group unless otherwise specified. This is based on the empirical results in Figure 6 that maintaining the embeddings of overlapping tokens close to the original status during the adaptation is effective only for non-Latin script languages and counter-effective for Latin languages. More details regarding the training are in Appendix A.
Full-Weight Fine-tuning.
After the adaptation phase, we follow the fine-tuning recipe of ALMA (Xu et al., 2024) that consists of full-weight training with monolingual corpus and a small amount of high-quality parallel corpus to enhance MT performance (§3.4). We also include Mistral in this phase of training. 1) In monolingual fine-tuning, we use MADLAD-400. For adapted ZeTT and FOCUS models, we fine-tune each separate model with non-English language-specific vocabulary except for uk and kk for ZeTT (again, due to unsupported languages in ZeTT) using a total of 2B tokens combining English and the corresponding non-English. For Mistral and the adapted VocADT and OFA, we fine-tune separate models for all three non-English groups (Latin, Mixed, Cyrillic) plus English using a corpus of 5B monolingual tokens containing 5 languages. 2) In the next parallel training, we sample 15k bitext from NLLB dataset [^5] https://huggingface.co/datasets/allenai/nllb for each English and non-English training pairs with top LASER3 scores (Artetxe & Schwenk, 2019; Schwenk et al., 2021; Heffernan et al., 2022; NLLB Team et al., 2022). The parallel training is done for one epoch, and we report test set numbers with the best model of the validation set. All the models are fine-tuned and tested with both directions of en-xx and xx-en within a single model, meaning there are no separate models for opposite translation directions. We follow the prompting strategy of Xu et al. (2024).

### v2

@@ Abstract
Vocabulary adaptation, which integrates new vocabulary into pre-trained language models, enables expansion to new languages and mitigates token over-fragmentation. However, existing approaches are limited by their reliance on heuristics or external embeddings. We propose VocADT, a novel method for vocabulary adaptation using adapter modules that are trained to learn the optimal linear combination of existing embeddings while keeping the model’s weights fixed. VocADT offers a flexible and scalable solution without depending on external resources or language constraints. Across 11 languages—with diverse scripts, resource availability, and fragmentation—we demonstrate that VocADT outperforms the original Mistral model (Jiang et al., 2023) and other baselines across various multilingual tasks including natural language understanding and machine translation. We find that Latin-script languages and highly fragmented languages benefit the most from vocabulary adaptation. We further fine-tune the adapted model on the generative task of machine translation and find that vocabulary adaptation is still beneficial after fine-tuning and that VocADT is the most effective. [^3] Project page: https://github.com/h-j-han/VocADT. Models at Huggingface Hub
@@ 1 Introduction
Vocabulary adaptation (or transfer)—a process of modifying a pre-trained language model (LM) to use a new vocabulary—offers several key advantages. First, it enables the introduction of new languages into a model, increasing flexibility in handling linguistic diversity and improving downstream performance in target languages (Wang et al., 2020; Gogoulou et al., 2022; Downey et al., 2023). Second, it reduces over-fragmentation, where words are excessively split by the tokenizer, slowing down generation [^4] Standard transformer decoding is quadratic in sequence length, so length increases can be catastrophic. and impairing performance in certain languages (Ahia et al., 2023; Petrov et al., 2023; Yamaguchi et al., 2024). These benefits have led to the development of numerous vocabulary adaptation approaches that initialize the new embeddings of new vocabulary with various methods based on heuristics (Mosin et al., 2023; Gee et al., 2022; Downey et al., 2023), external resources (Tran, 2020; Dobler & de Melo, 2023; Liu et al., 2024), or a separate hypernetwork that generates it (Minixhofer et al., 2024). They generally generate new embeddings using original embeddings, optionally followed by continued training to finalize the adaptation (Minixhofer et al., 2022; Ostendorff & Rehm, 2023; Dobler & de Melo, 2023; Liu et al., 2024).
@@
However, existing vocabulary adaptation approaches face several limitations. Those that rely on heuristics (Gee et al., 2022; Downey et al., 2023), which use predefined rules to initialize new embeddings from existing ones rather than learning from data, often lack adaptability, where the new embeddings are not fully integrated into the original model and require an additional training phase of full-weight updates to fully adapt to the new vocabulary. Also, those that depend on external embeddings or networks (Tran, 2020; Dobler & de Melo, 2023; Liu et al., 2024), increase complexity and limit scalability. Furthermore, some approaches focus solely on language-specific cases or may have restrictions on the number of languages in the implementation when configuring the vocabulary (Dobler & de Melo, 2023; Minixhofer et al., 2024).
@@
Additionally, we still know little about the impact of vocabulary adaptation across diverse linguistic and task settings. Most prior work investigates few languages, which is insufficient for identifying patterns (Tran, 2020; Ostendorff & Rehm, 2023; Remy et al., 2023; Yamaguchi et al., 2024), while studies that consider a broader range of languages only report averages without detailed analysis (Liu et al., 2024; Mundra et al., 2024). Furthermore, the impact of vocabulary adaptation on cross-lingual and generative tasks like machine translation (MT) is understudied, even though they represent crucial application areas for porting models to new languages. Many adaptation methods (Chung et al., 2020; Gee et al., 2022; Liu et al., 2024) have been evaluated instead on non-cross-lingual and discriminative tasks such as commonsense reasoning, natural language inference (NLI), or question answering (QA)—which are typically classification tasks.
@@
We propose VocADT, a novel solution for vocabulary adaptation using adapters, designed to address key challenges in existing approaches (Figure 1). We introduce a vocabulary adapter module, a learnable matrix between the new vocabulary and the original embeddings of a language model. The module gradually adapts to new vocabularies through training while keeping all weights of the original model fixed, allowing the module to learn the best combination of the original embeddings without relying on heuristics, external embeddings, or dictionaries. This learned adaptation approach offers better adaptability of new embeddings to the original language model with only its embeddings replaced and more flexibility in the number of languages while removing the necessity of external pre-trained resources. At the end of training, the adapter is merged with the original embeddings to create a new embedding matrix.
@@
In addition to our novel method, we empirically address the following key questions to understand the effectiveness and behavior of vocabulary adaptation: (1) Which languages benefit most from vocabulary adaptation? (2) What are the best strategies for creating new vocabularies? Also, is script consistency necessary? (3) How does vocabulary adaptation impact machine translation? We emphasize this task as it is a critical task for multilingual models that involves cross-lingual and generative capabilities, which are often more complex than classification or monolingual tasks.
@@
Figure 1: Overview of our vocabulary adaptation with adapter (VocADT) and the initialization of adapter. The vocabulary adapter modules are trained to adapt new vocabulary with existing embeddings while keeping the original model fixed. We initialize entries of the adapter for overlapping tokens and tokens whose partitions are in the original vocabulary. Once trained, the adapters and original embeddings are merged to form the new embeddings.
We demonstrate the effectiveness of our adaptation method on various NLP tasks spanning Natural Language Understanding and MT. Results show that our approach consistently surpasses the original Mistral model in most cases, both after the adaptation phase and following phase of full-weight training. Additionally, our method outperforms or matches other strong vocabulary adaptation baselines. Our findings indicate that Latin-script languages and those with severe fragmentation benefit the most from vocabulary adaptation. Finally, while all vocabulary adaptation methods continue to be effective for machine translation after fine-tuning, VocADT shows the best results among them. Our main contributions are summarized as follows:
  • We propose VocADT, a simple and effective solution for vocabulary adaptation using adapters, that addresses key limitations in prior work such as reliance on external embedding or language constraints.
  • We conduct experiments that cover a wide range of languages and scripts, finding that languages with Latin scripts or severe fragmentation benefit the most and that having a consistent grouping of scripts for multilingual vocabulary is helpful.
  • Our approach consistently outperforms the original language model and is more effective than, or on par with, strong vocabulary adaptation baselines after the adaptation phase across various tasks and after the following full-weight fine-tuning on MT.
2 Background and Motivation
Approaches to Vocabulary Adaptation
Prior work focuses on initialization strategies for the new vocabulary embeddings, before continuing training with unlabeled target language text using the original self-supervised pretraining objective. For instance, FOCUS (Dobler & de Melo, 2023) initializes embeddings as a weighted combination of overlapping tokens using external embeddings for non-overlapping ones, while copying the embeddings of overlapping tokens. OFA (Liu et al., 2024) also relies on external word vectors to initialize embeddings for non-shared new tokens, using a weighted average of original tokens based on semantic similarity. This strategy often requires external resources such as auxiliary embeddings (Dobler & de Melo, 2023; Liu et al., 2024; Ostendorff & Rehm, 2023) or bilingual dictionaries (Mundra et al., 2024; Minixhofer et al., 2022).
After initialization, language adaptive pretraining (LAPT; Chau et al., 2020) usually updates all model weights (Tran, 2020; Liu et al., 2024; Dobler & de Melo, 2023; Ostendorff & Rehm, 2023), except Yamaguchi et al. (2024) which use LoRA (Hu et al., 2022). Downey et al. (2023) show full-weight updates outperform embedding-only training, which is insufficient for multilingual transfer.
@@ 2 Background and Motivation
Linear Combination of Embeddings
Most vocabulary transfer methods combine the existing embeddings to generate new ones. A popular approach is to use a weighted average of the original embeddings (bolded in Appendix Table 9). For example, Gee et al. (2022) and Mosin et al. (2023) compute the new embeddings by simply averaging the embeddings of subword tokens, while Tran (2020), Minixhofer et al. (2022), OFA, and FOCUS utilize external resources to determine the weights to initialize the new embeddings with a weighted average of the original embeddings. Mundra et al. (2024) established theoretically that initializing within the convex hull of existing embeddings—e.g., using a weighted average of source embeddings—is a good initialization.
Our motivation stems from the question: rather than deciding how to combine existing embedding vectors heuristically, why not learn this process to create new embedding vectors? Relying on heuristics may lack adaptability that typically requires an additional training phase of full-weight updates to fully adapt to the new vocabulary. Building upon prior works, we propose to learn linear combinations with vocabulary adapters.
@@
Furthermore, most studies exclusively evaluate models on non-cross-lingual and non-generative tasks, such as binary or multi-class classification, sequence labeling (e.g., part-of-speech tagging), or answer span prediction. Mundra et al. (2024) and Yamaguchi et al. (2024) include monolingual generative tasks like summarization. As a result, the impact of vocabulary adaption on cross-lingual generation tasks such as MT remains understudied, even though this is a crucial application area. 555 Mundra et al. (2024) report MT results, but they do not release the models or per-language performance metrics, making direct comparisons difficult.
To address these gaps, this work introduces a strategy to adapt LLM to new vocabularies, and experiments designed to measure its impact across diverse linguistic and task settings. We compare empirically against FOCUS and OFA, representing more resource-intensive initialization approaches, and ZeTT, representing a more parsimonious approach that has only been tested on limited tasks so far. Summaries of prior work can be found in Appendix Table 9.
@@ 3.1 Vocabulary Adapter Module
We introduce the vocabulary adapter modules to find parameters of new embeddings that can replace the original embedding without changing the non-embedding part of the original model (Figure 1(a)). For simplicity, we refer to both input and output embeddings (or the LM head) collectively as embeddings. Let $V^{o}$ and $V^{n}$ be the $\bm{o}$riginal and $\bm{n}$ew vocabulary, respectively, and let $\mathcal{T}^{x}:w\rightarrow(t_{1},t_{2},\dots,t_{k})$ be a tokenizer associated with a vocabulary $V^{x}$ where $t_{j}\in V^{x},\forall j=1,\dots,k$. We put vocabulary adapter modules ${\bm{A}}\in\mathbb{R}^{|V^{n}|\times|V^{o}|}$ between the new vocabulary $V^{n}$ and the original embedding ${\bm{E}}^{o}\in\mathbb{R}^{|V^{o}|\times h}$ where $h$ is an embedding dimension, in a manner similar to bottleneck adapters (Houlsby et al., 2019). We train the adapters with the standard language modeling loss $\mathcal{L}^{lm}$, where we freeze the original weights and only update the adapters. This may be analogous to finding the new embedding vector for a token with the weighted combination of original embedding vectors (Downey et al., 2023; Dobler & de Melo, 2023; Liu et al., 2024). Unlike similar works, our approach learns the weights for embedding combination. After training the adapters, we get new embeddings ${\bm{E}}^{n}\in\mathbb{R}^{|V^{n}|\times h}$ by merging the original embeddings and adapters to ${\bm{E}}^{n}={\bm{A}}{\bm{E}}^{o}$, which result in a language model with the same architecture as the original one but with a different vocabulary size.
@@ 3.4 Further Fine-tuning For Downstream Task
To understand the impact of vocabulary adaptation after task-specific fine-tuning, we follow the full ALMA (Xu et al., 2024) training on all model parameters for the cross-lingual generation task of machine translation after our VocADT on just the embeddings. ALMA training begins with fine-tuning on monolingual data, followed by further weight optimization on small curated parallel data.
@@ 4 Which Languages Benefit the Most from Vocabulary Adaptation?
We aim to understand “When and how should we perform vocabulary adaptation?”. More specifically, we seek insight into which languages might benefit the most from vocabulary adaptation in terms of improving overall performance or mitigating over-fragmentation.666 The non-English languages that we cover are all highly fragmented by common LLMs, and their fragmentation is simmilarly improved by our method. Therefore, our analysis focuses on performance.
To this end, we design experiments to cover 10 non-English languages along with English, listed in Table 1, with a variety of scripts and language families. These languages are broadly categorized into three groups: (1) Latin group of Swahili, Indonesian, Estonian, and Haitian, which are low- to mid-resource languages and all use Latin script; (2) Mixed group including Korean, Greek, Russian, and Bulgarian, which utilize a mixture of scripts; (3) Cyrillic group for languages with that scrip. [^7] In Section 6.4 and Appendix F.1, we experiment with All group including all languages mentioned here.
@@ 5.2 Training VocADT
In the adapter training phase, we train only the adapters, while fixing all parameters of the original model. The input and output adapters are separate modules, as preliminary results showed that sharing an adapter for the input and output sides performs worse. We train 0.5B monolingual tokens per language, totaling 2.5B mixed by 5 languages (English + 4 non-English from each corresponding group), and report test numbers from it. We use “clean” documents from the corpus of MADLAD-400 (Kudugunta et al., 2023). We set the weighing factor of auxiliary loss $\alpha$ with 0.1 for non-Latin groups and 0 for the Latin group unless otherwise specified. This is based on the empirical results in Appendix A that maintaining the embeddings of overlapping tokens close to the original status during the adaptation is effective only for non-Latin script languages and counter-effective for Latin languages. More details regarding the training are in Appendix C.
5.3 Full-Weight Fine-tuning
After the adaptation phase, we follow the fine-tuning recipe of ALMA (Xu et al., 2024) that consists of full-weight training with monolingual corpus and a small amount of high-quality parallel corpus to enhance MT performance (§3.4). We also include Mistral in this phase of training. 1) In monolingual fine-tuning, we use MADLAD-400. For adapted ZeTT and FOCUS models (prior work), we fine-tune each separate model with non-English language-specific vocabulary except for uk and kk for ZeTT (again, due to unsupported languages in ZeTT) using a total of 2B tokens combining English and the corresponding non-English. For Mistral and the adapted VocADT and OFA, we fine-tune separate models for all three non-English groups (Latin, Mixed, Cyrillic) plus English using a corpus of 5B monolingual tokens containing 5 languages. 2) In the next parallel training, we sample 15k bitext from NLLB dataset (Schwenk et al., 2021b; Heffernan et al., 2022; NLLB Team et al., 2022) [^9] https://huggingface.co/datasets/allenai/nllb for each English and non-English training pairs with top LASER3 scores (Artetxe & Schwenk, 2019). The parallel training is done for one epoch, and we report test set numbers with the best model of the validation set. All the models are fine-tuned and tested with both directions of en-xx and xx-en within a single model, meaning there are no separate models for opposite translation directions. We follow the prompting strategy of Xu et al. (2024).
Figure 2: Average scores of original Mistral and its adaptation with new vocabulary, only replacing embeddings and fixing the body of transformer modules. “-multi” indicates models with a multilingual vocabulary, which includes five languages covering all languages with two separate models, while “-mono” refers to monolingual vocabulary models. xx-en and en-xx indicate MT tasks. See Appendix E for individual values.